# Variational Gaussian Approximation in Replica Analysis of Parametric Models

### v1

@@ Appendix A Grand canonical
$\frac{\sum_{k=0}^{\infty}\frac{e^{-n}n^{k}}{k!}\mathbb{E}_{z\sim p_{\rm{data}}}\left[e^{-\beta\sum_{a=1}^{r}l(\theta^{a};z)}\right]^{k}e^{-\frac{\beta\lambda}{2}\sum_{a=1}^{r}\|w^{a}\|_{2}^{2}}}{\sum_{k=0}^{\infty}\frac{e^{-n}n^{k}}{k!}\,Z_{r}^{\beta,k}}$
@@
This expression coincides with the GC formalism  (13). From the normalization it is evident that this formulation is formally similar to a grand canonical ensemble of ideal gas in statistical mechanics, with the sample size playing the role of a fluctuating particle number and $\beta^{-1}\log n$ acting as the chemical potential. In this sense, $\tilde{D}$ may be regarded as a reservoir. This analogy provides an intuitive physical interpretation of the GC formalism. It would also be useful when considering more elaborate replica systems, such as those used in the replica analysis of bootstrap methods Takahashi (2023, 2024, 2025).

### v2

@@ Appendix A Grand canonical
$\frac{\sum_{k=0}^{\infty}\frac{e^{-n}n^{k}}{k!}\mathbb{E}_{z\sim p_{\rm{data}}}\left[e^{-\beta\sum_{a=1}^{r}l(\theta^{a};z)}\right]^{k}e^{-\frac{\beta\lambda}{2}\sum_{a=1}^{r}\|w^{a}\|_{2}^{2}}}{\sum_{k=0}^{\infty}\frac{e^{-n}n^{k}}{k!}\Xi_{r}^{\beta,k}}$
@@
This expression coincides with the GC formalism (13).
The representation of the normalization constant implies that in GC formalism, the partition function can be written as
$\Xi_{r,{\rm GC}}^{\beta,n}=\sum_{k=0}^{\infty}n^{k}\frac{1}{k!}\Xi_{r}^{\beta,k},$ (86)
This transform is similar to the one from canonical ensemble to grand canonical ensemble with fugacity $n$. In this sense, $\tilde{D}$ can be regarded as a reservoir. Also, it may be possible to interpret that the factor $k!$ reflect the label-exchange symmetry of the data points. The computation in this section is equivalent to the one in the main text. Nevertheless, we expect that it would be useful when considering more elaborate replica systems, such as those used in the replica analysis of bootstrap methods Malzahn and Opper (2003b); Obuchi and Kabashima (2019); Takahashi (2023, 2024, 2025).